How far must a 1.00x10-7C test charge (q) be from a 5.00x10-6C point

charge (Q) for it to experience a push of about 1.25 N?


1
Expert's answer
2022-05-22T14:39:19-0400
"F=k\frac{q_1q_2}{r^2}"

"r=\sqrt{\frac{kq_1q_2}{F}}\\
=\sqrt{\frac{9*10^9*1.00*10^{-7}*5.00*10^{-6}}{1.25}}=0.06\:\rm m=6\: cm"